# High-dimensional Gaussian and bootstrap approximations for robust means

### v3

@@ 1 Introduction
Let $X_{1},\ldots,X_{n}$ be a sample of i.i.d. random vectors in $\mathbb{R}^{d}$ with mean vector $\mu$ and covariance matrix $\Sigma$. Furthermore, let $S_{n}=n^{-1/2}\sum_{i=1}^{n}\mathinner{(X_{i}-\mu)}$. Since the seminal paper of Chernozhukov et al. (2013) there has been substantial interest in Gaussian approximations to the distribution of $S_{n}$ when $d$ is large relative to $n$. Letting $Z\sim\mathsf{N}_{d}(0,\Sigma)$ and $\mathcal{H}$ be the class of (generalized) hyperrectangles in $\mathbb{R}^{d}$, that is the class of all sets of the form
@@
and related quantities, cf., e.g., Chernozhukov et al. (2017a); Deng and Zhang (2020); Lopes et al. (2020); Kuchibhotla and Rinaldo (2020); Das and Lahiri (2021); Koike (2021); Kuchibhotla et al. (2021); Lopes (2022); Chernozhuokov et al. (2022); Fang et al. (2023); Chernozhukov et al. (2023b); Koike (2024). We refer to the review in Chernozhukov et al. (2023a) for further references. For example, when the entries of $X_{i}=(X_{i,1},\ldots,X_{i,d})^{\prime}$ are (uniformly) sub-exponential, $\rho_{n}\to 0$ if $d=d(n)=o\mathinner{\bigl(\exp(n^{1/5})\bigr)}$, cf. Chernozhuokov et al. (2022). Thus, $d$ can grow exponentially fast with $n$ and this rate can be further improved under additional assumptions on the distribution of the $X_{i}$ such as, e.g., variance decay conditions on $\Sigma$ as in Lopes et al. (2020) or eigenvalue conditions as in Fang and Koike (2021); Kuchibhotla and Rinaldo (2020); Chernozhukov et al. (2023b).
@@
Despite the progress on such high-dimensional Gaussian approximations for $S_{n}$, it follows from Remark 2 in Zhang and Wu (2017) and Theorem 2.1 in Kock and Preinerstorfer (2024) that for every $m\in(2,\infty)$ there exist i.i.d. random vectors $X_{1},\ldots,X_{n}$ with independent entries $X_{ij}\sim P_{m}$, and $P_{m}$ depending neither on $n$ nor $d$, having mean zero, variance one, and finite $m$th absolute moment, such that if for some $\xi\in(0,\infty)$ it holds that $\limsup_{n\to\infty}\frac{d}{n^{m/2-1+\xi}}>0$, then $\limsup_{n\to\infty}\rho_{n}=1$. In particular, for any given $m\in(2,\infty)$, the Gaussian approximation $\rho_{n}\to 0$ does not hold uniformly over all distributions with bounded $m$th moments when $d$ grows exponentially in $n$.
@@
Conversely, it is a simple consequence of Theorem 2 in Chernozhukov et al. (2023a), cf. Theorem 2.2 in Kock and Preinerstorfer (2024), that $\rho_{n}\to 0$ uniformly over a large class of distributions with bounded $m$th moments if there exists a $\xi\in(0,\infty)$ such that $\lim_{n\to\infty}\frac{d}{n^{m/2-1-\xi}}=0$. Hence, a critical phase transition occurs for the asymptotic behaviour of $\rho_{n}$ at $d=n^{m/2-1}$. As $d$ passes this threshold from below, the limit of $\rho_{n}$ jumps from zero to one. For example, for $m=3$ one can construct $X_{i,j}$ with bounded third moments such that $\rho_{n}\to 1$ if $d=n^{1/2+\xi}$ for $\xi$ arbitrarily close to zero. Thus, even in a regime where $d$ grows (much) slower than $n$, the Gaussian approximation to the distribution of $S_{n}$ can break down completely if the $X_{i}$ only possess three moments.
@@
Motivated by this phase transition, Resende (2024) recently studied the case where $S_{n}$ is replaced by a suitably trimmed mean and $\mathcal{H}$ is replaced by the subfamily of “one-sided” intervals $\mathcal{R}$, say, i.e., the class of all sets of the form $R=\mathinner{\{x\in\mathbb{R}^{d}\mathrel{\mathop{\ordinarycolon}}x_{j}\leq t_{j}\text{ for all }j=1,\ldots,d\}}$, where $t_{j}\in\mathbb{R}$ for all $j=1,\ldots,d$. He obtained Gaussian approximations that are informative even when the $X_{i}$ only possess $m>2$ moments and $d$ grows exponentially fast in $n$. This is of fundamental importance, as it shows that one can break through the barrier $d=n^{m/2-1}$ faced by $\rho_{n}$, which is based on $S_{n}$. The exact permitted growth rate of $d$ depends on $m$. As $m\to\infty$, his result allows $d$ to grow almost as fast as $\exp(n^{1/6})$ for the Gaussian approximation as well as an empirical bootstrap and as fast as $\exp(n^{1/8})$ for a multiplier bootstrap. Furthermore, the trimming ensures that these approximations remain valid even when some of the $X_{i}$ have been adversarially contaminated prior to being given to the statistician. This is in stark contrast to statistics based on the sample mean $S_{n}$, which have a breakdown point of $1/n$ ($S_{n}$ can be changed to any value by manipulating only one of the vectors $X_{i}$). A potential drawback of the trimmed mean studied in Resende (2024) is that the amount of trimming needed depends on $m$ (cf. Theorem 2 in Resende (2024)), which is typically unknown in practice. Thus, if one constructs an estimator based on an $m$ higher than the actual number of moments that the $X_{i}$ possess one does not have any approximations guarantees for the trimmed mean, whereas the guarantees one obtains may be suboptimal if the $X_{i}$ possess more moments than used in the construction of the trimmed mean.
@@
We also mention the work of Liu and Lopes (2024) who, motivated by the poor performance of $S_{n}$ in the presence of heavy tails, even established a dimension-independent bootstrap approximation over $\mathcal{R}$ for certain robust max statistics related to the winsorized means we study under the conditions of $L^{4}$-$L^{2}$ moment equivalence, a variance decay condition on $\Sigma$, and restrictions on the Frobenius norm of certain submatrices of the correlation matrix of the $X_{i}$. Robustness to outliers or other sources of (adversarial) data contamination were not investigated.
In this paper we do not impose any structural assumptions on $\Sigma$ (apart from positive variances) and obtain Gaussian approximations to the distributions of winsorized and trimmed means over $\mathcal{H}$, which contains the family $\mathcal{R}$ studied in Resende (2024). The winsorization and trimming points are suitably chosen order statistics and we only require $\log(d)=o\mathinner{\bigl(n^{\frac{m-2}{5m-2}}\bigr)}$, which is exponential for all $m>2$ albeit with a small exponent for $m$ close to two. Apart from our Gaussian approximations being valid over a larger family of sets, an important advantage over the trimmed mean studied in Resende (2024) is that one does not need to know the number of moments $m$ that the $X_{i}$ possess in order to implement our winsorized and trimmed means — they “adapt” to $m$. Furthermore, as $m\to\infty$, we allow $d$ to grow almost as fast as $\exp(n^{1/5})$, improving on the rate in Resende (2024), and thus “recover” the best known rate (cf. Remark 2 in Chernozhukov et al. (2023a)) for Gaussian approximations based on the sample mean of $X_{i}$ with sub-exponential entries. This rate remains valid for the bootstrap procedures that we consider and all results are robust to some adversarial contamination. Our bootstrap approximations are based on a novel covariance matrix estimator, for which we establish performance guarantees in the presence of adversarial contamination and only $m>2$ moments in Section 3.1. This estimator does not require knowledge of any unknown population quantities. In particular, its performance guarantees adapt to the unknown $m$, which may be of independent interest.
@@
In contrast to the present paper, which focuses exclusively on the canonical problem of Gaussian approximations in $\mathbb{R}^{d}$Resende (2024) also considers Gaussian approximations over VC-subgraph classes of functions and applies his results to vector mean estimation under general norms.
@@ 2 Gaussian approximations for winsorized means
Recall that $X_{1},\ldots,X_{n}$ is a sample of i.i.d. random vectors in $\mathbb{R}^{d}$ with $X_{i}=(X_{i,1},\ldots,X_{i,d})^{\prime}$ for $i=1,\ldots,n$. Let $\mu=(\mu_{1},\ldots,\mu_{d})^{\prime}=\mathbb{E}X_{1}$, $\Sigma$ be the covariance matrix of $X_{1}$, and for $m\in[2,\infty)$ let $\sigma_{m,j}^{m}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}|X_{1,j}-\mu_{j}|^{m}$, all of which are well-defined under Assumption 2.1 below. We suppress the dependence of $d=d(n)$ on $n$ in our notation.
@@
where $\overline{\eta}_{n}\in(0,1/2)$ is a non-random and known upper bound on the fraction of contaminated observations. Which of $\tilde{X}_{i}$ differ from $X_{i}$ as well as their values can depend on the uncontaminated sample $X_{1},\ldots,X_{n}$. Adversarial contamination has become a popular criterion to evaluate robustness of a statistic against as it allows for many forms of data manipulation, cf. Lai et al. (2016), Cheng et al. (2019), Diakonikolas et al. (2019), Hopkins et al. (2020), Lugosi and Mendelson (2021), Minsker and Ndaoud (2021), Bhatt et al. (2022), Depersin and Lecué (2022), Dalalyan and Minasyan (2022), Minasyan and Zhivotovskiy (2023), Minsker (2023), Oliveira et al. (2025). The recent book by Diakonikolas and Kane (2023) provides further references and discussion of various contamination settings. Since the sample mean has a breakdown point of $1/n$, Gaussian approximations based on $S_{n}$ are not robust to adversarial contamination (or large outliers).
@@
The $X_{1},\ldots,X_{n}$ are i.i.d. random vectors in $\mathbb{R}^{d}$ with, $\mathbb{E}|X_{1,j}|^{m}<\infty$ for some $m\in(2,\infty)$ and all $j=1,\ldots,d$. Suppose that there exist strictly positive constants $b_{1}$ and $b_{2}$ such that $\min_{j=1,\ldots,d}\sigma_{2,j}\geq b_{1}$ and $\sigma_{m}\mathrel{\mathop{\ordinarycolon}}=\max_{j=1,\ldots,d}\sigma_{m,j}\leq b_{2}$. The actually observed adversarially contaminated random vectors (in $\mathbb{R}^{d}$) are denoted $\tilde{X}_{1},\ldots,\tilde{X}_{n}$ and satisfy (2).
@@ 2.1 The winsorized means
with $\hat{\alpha}_{j}=\tilde{X}_{\lceil\varepsilon_{n}n\rceil,j}^{*}$ and $\hat{\beta}_{j}=\tilde{X}_{\lceil(1-\varepsilon_{n})n\rceil,j}^{*}$ for $\varepsilon_{n}\in(0,1/2)$. Thus, for each coordinate $j$, the winsorization points $\hat{\alpha}_{j}$ and $\hat{\beta}_{j}$ are order statistics of the contaminated data $\tilde{X}_{1,j},\ldots,\tilde{X}_{n,j}$. [^2] For the purpose of construction estimators of $\mu\in\mathbb{R}$ with finite-sample sub-Gaussian concentration properties, related winsorized mean estimators were recently studied in Lugosi and Mendelson (2021) and Kock and Preinerstorfer (2025). Under adversarial contamination it is clear that even $S_{n,W}$ can perform arbitrarily badly unless at least the smallest and largest $\overline{\eta}_{n}n$ observations are winsorized. Thus, one must choose $\varepsilon_{n}\geq\overline{\eta}_{n}$. In particular, we study $\varepsilon_{n}$ of the form
$\varepsilon_{n}=\lambda_{1}\cdot\overline{\eta}_{n}+\lambda_{2}\cdot\frac{\log(dn)}{n},\qquad\lambda_{1}\in(1,\infty)\text{ and }\lambda_{2}\in(0,\infty).$
To make efficient use of the data, it is desirable to establish Gaussian approximations with $\varepsilon_{n}$, and thus $\lambda_{1}$ and $\lambda_{2}$, as small as possible. We now discuss the choice of $\lambda_{1}$ and $\lambda_{2}$.
Consider first the case of $\overline{\eta}_{n}=0$ (no contamination), which is the setting in which high-dimensional Gaussian approximations for the sample mean based $S_{n}$ have been studied primarily (cf. the literature summarized in the introduction). In this case, one can choose $\lambda_{2}=6.05$ whereas the choice of $\lambda_{1}$ is irrelevant. In fact, as will be seen in Remark 2.1, even smaller choices of $\lambda_{2}$ are possible.
When one suspects that the data may have been contaminated, corresponding to $\overline{\eta}_{n}>0$, there is a tradeoff between the sizes of $\lambda_{1}$ and $\lambda_{2}$ in our implementation of the winsorized means, which we parameterize by $c\in(1,\sqrt{1.5})$. In particular, following Kock and Preinerstorfer (2025), one can choose
$\lambda_{1}=\lambda_{1,c}=\frac{c}{1-\sqrt{2(c^{2}-1)}}$
and
$\lambda_{2}=\lambda_{2,c}=\lambda_{2,c}(n,d)=\mathinner{\biggl[\frac{c}{3[1-\sqrt{2(c^{2}-1)}]}\vee c\mathinner{\biggl(\sqrt{2\frac{c+1}{c-1}}+\frac{1}{3}\biggr)}\biggr]}\wedge c\mathinner{\biggl(\sqrt{\frac{n}{2\log(dn)}}+\frac{1}{3}\biggr)}.$ (4)
In the sequel $\varepsilon_{n}$ refers to (we suppress its dependence on $c\in(1,\sqrt{1.5})$ and $d$)
$\varepsilon_{n}=\lambda_{1,c}\cdot\overline{\eta}_{n}+\lambda_{2,c}\cdot\frac{\log(dn)}{n}.$ (5)
Since $c\mapsto\lambda_{1,c}$ is a (strictly increasing) bijection from $(1,\sqrt{1.5})$ to $(1,\infty)$, any value of $\lambda_{1,c}\in(1,\infty)$ can be achieved by a suitable choice of cc. [^3] To achieve $\lambda_{1,c}=A\in(1,\infty)$, set $c=\frac{\sqrt{2}\sqrt{3A^{4}-A^{2}}-A}{2A^{2}-1}$.
Remark 2.1.
In the important case of $\overline{\eta}_{n}=0$ one can nearly minimize $\lambda_{2,c}$, and thus the number of winsorized observations, by checking whether i) $c=\tilde{c}$ with $\tilde{c}$ being the minimizer of the term $f(c)$ in square brackets in (4) or ii) $c$ arbitrarily close to one yields the smallest value of $\lambda_{2,c}$. [^4] To see this, note that $f(c)=\mathinner{\Bigl[\frac{c}{3[1-\sqrt{2(c^{2}-1)}]}\vee c\mathinner{\bigl(\sqrt{2\frac{c+1}{c-1}}+\frac{1}{3}\bigr)}\Bigr]}$ is minimized by equating the two terms at $\tilde{c}=\frac{1}{17}(-4+3\sqrt{66})\approx 1.198$ which results in $f(\tilde{c})=6.041346$ whereas $c\mapsto c\mathinner{\Bigl(\sqrt{\frac{n}{2\log(dn)}}+\frac{1}{3}\Bigr)}$ is strictly increasing in $c$ for given values of $n$ and $d$. In particular, $\lambda_{2,c}$ chosen in this way will never exceed $6.05$ since $f(\tilde{c})\leq 6.05$.

### v4

@@ 1 Introduction
Let $X_{1},\ldots,X_{n}$ be a sample of i.i.d. random vectors in $\mathbb{R}^{d}$ with mean vector $\mu$ and covariance matrix $\Sigma$. Furthermore, let $S_{n}=n^{-1/2}\sum_{i=1}^{n}\mathinner{(X_{i}-\mu)}$. Since the seminal paper of Chernozhukov et al. (2013) there has been much renewed interest in Gaussian approximations to the distribution of $S_{n}$ when $d=d(n)$ is large relative to $n$. Letting $Z\sim\mathsf{N}_{d}(0,\Sigma)$ and $\mathcal{H}$ be the class of (generalized) hyperrectangles in $\mathbb{R}^{d}$, that is the class of all sets of the form
@@
and related quantities, cf., e.g., Chernozhukov et al. (2017a); Deng and Zhang (2020); Lopes et al. (2020); Kuchibhotla and Rinaldo (2020); Das and Lahiri (2021); Koike (2021); Kuchibhotla et al. (2021); Lopes (2022); Chernozhukov et al. (2022); Fang et al. (2023); Chernozhukov et al. (2023b); Koike (2024). We refer to the review in Chernozhukov et al. (2023a) for further references. For example, when the entries of $X_{i}=(X_{i,1},\ldots,X_{i,d})^{\prime}$ are (uniformly) sub-exponential, $\rho_{n}\to 0$ if $\log(d)=o\mathinner{\bigl(n^{1/5}\bigr)}$, cf. Chernozhukov et al. (2022). Thus, $d$ can grow exponentially fast with $n$ and this rate can be further improved under additional assumptions on the distribution of the $X_{i}$ such as, e.g., variance decay conditions on $\Sigma$ as in Lopes et al. (2020) or eigenvalue conditions as in Fang and Koike (2021); Kuchibhotla and Rinaldo (2020); Chernozhukov et al. (2023b).
@@
Despite the progress on such high-dimensional Gaussian approximations for $S_{n}$, it follows from Remark 2 in Zhang and Wu (2017) and Theorem 2.1 in Kock and Preinerstorfer (2024) that for every $m\in(2,\infty)$ there exist i.i.d. random vectors $X_{1},\ldots,X_{n}$ with independent entries $X_{i,j}\sim P_{m}$, and $P_{m}$ depending neither on $n$ nor $d$, having mean zero, variance one, and finite $m$th absolute moment, such that if for some $\xi\in(0,\infty)$ it holds that $\limsup_{n\to\infty}\frac{d}{n^{m/2-1+\xi}}>0$, then $\limsup_{n\to\infty}\rho_{n}=1$. In particular, for any given $m\in(2,\infty)$, the Gaussian approximation $\rho_{n}\to 0$ does not hold uniformly over all distributions with bounded $m$th moments when $d$ grows exponentially in $n$. Figure 1 illustrates how the Gaussian approximation to $S_{n}$ can fail severely over $\mathcal{H}$ when the $X_{i,j}$ have relatively heavy tails. In particular, letting $x||_{\infty}=\max_{1\leq j\leq d}|x_{j}|$ for $x\in\mathbb{R}^{d}$, it is seen that $S_{n}||_{\infty}$ has much heavier tails than $Z||_{\infty}$.
@@
Conversely, it is a simple consequence of Theorem 2 in Chernozhukov et al. (2023a), cf. Theorem 2.2 in Kock and Preinerstorfer (2024), that $\rho_{n}\to 0$ uniformly over a large class of distributions with bounded $m$th moments if there exists a $\xi\in(0,\infty)$ such that $\lim_{n\to\infty}\frac{d}{n^{m/2-1-\xi}}=0$. Hence, a critical phase transition occurs for the asymptotic behaviour of $\rho_{n}$ at $d=n^{m/2-1}$. As $d$ passes this threshold from below, the limit of $\rho_{n}$ jumps from zero to one.
@@
Motivated by this phase transition, Resende (2024) recently studied the case where $S_{n}$ is replaced by a suitably trimmed mean and $\mathcal{H}$ is replaced by the subfamily $\mathcal{R}$ consisting of all sets of the form $R=\mathinner{\{x\in\mathbb{R}^{d}\mathrel{\mathop{\ordinarycolon}}x_{j}\leq t\text{ for all }j=1,\ldots,d\}}$, where $t\in\mathbb{R}$. He obtained Gaussian approximations that are informative even when the $X_{i}$ only possess $m>2$ moments and $d$ grows exponentially fast in $n$. This is of fundamental importance, as it shows that one can break through the barrier $d=n^{m/2-1}$ faced by $\rho_{n}$, which is based on $S_{n}$. The exact permitted growth rate of $d$ depends on $m$. As $m\to\infty$, his result allows $d$ to grow almost as fast as $\exp(n^{1/6})$ for the Gaussian approximation as well as an empirical bootstrap and as fast as $\exp(n^{1/8})$ for a multiplier bootstrap. Furthermore, the trimming ensures that these approximations remain valid even when some of the $X_{i}$ have been adversarially contaminated prior to being given to the statistician. This is in stark contrast to statistics based on the sample mean $S_{n}$, which have a breakdown point of $1/n$ ($S_{n}$ can be changed to any value by manipulating only one of the vectors $X_{i}$).
@@
We also mention the work of Liu and Lopes (2024) who, motivated by the poor performance of $S_{n}$ in the presence of heavy tails, even established a dimension-independent bootstrap approximation over $\mathcal{R}$ for certain robust max-statistics related to the winsorized means we study (but working with a “variance-based” instead of a “quantile-based” winsorization) under the conditions of $L^{4}$-$L^{2}$ moment equivalence, a variance decay condition on $\Sigma$, and restrictions on the Frobenius norm of certain submatrices of the correlation matrix of the $X_{i}$. Robustness to outliers or other sources of (adversarial) data contamination were not investigated.
$n=200$ and d=5,000d=5{,}000               n=1,000n=1{,}000 and d=10,000d=10{,}000
[…]
@@
In the present paper, we obtain Gaussian approximations to the distributions of winsorized and trimmed means over $\mathcal{H}$ without imposing any structural assumptions on $\Sigma$ (apart from positive variances). Our Gaussian approximations are valid over all of $\mathcal{H}$ rather than the smaller family $\mathcal{R}$ studied in Resende (2024), which is crucial for the study of power properties of max-type tests based on our winsorized and trimmed means; cf. the results in the companion paper Kock and Preinerstorfer (2026). In contrast to the present paper, which focuses exclusively on the canonical problem of Gaussian approximations in $\mathbb{R}^{d}$Resende (2024) also considers Gaussian approximations over VC-subgraph classes of functions and applies his results to vector mean estimation under general norms.
One limitation of the trimmed mean studied in (Theorem 2) in Resende (2024) is that the amount of trimming needed depends on $m$, which is typically unknown in practice. If one implements the trimmed mean studied in Resende (2024) based on an $m$ higher than the actual number of moments that the $X_{i}$ possess, one does not have any approximation guarantees, whereas the guarantees one obtains may be suboptimal if the $X_{i}$ possess more moments than the $m$ used in the implementation of the trimmed mean. What is more, even when $m$ is known, it is often the case that the required number of trimmed observations for the estimator studied in Resende (2024)
  • i)
    either exceeds the sample size $n$, making that trimmed mean unimplementable (cf. the left panel of Figure 1 where the trimmed mean studied in Resende (2024) is absent);
  • ii)
    or is smaller than $n$ but so large that the Gaussian approximation to the distribution of that trimmed mean is not practically useful (cf. the right panel of Figure 1 showing that the tails of the trimmed mean are much lighter than those of the approximating Gaussian due to excessive trimming).
In contrast, to implement the winsorized and trimmed means we investigate, one does not need to know the number of moments $m$ that the $X_{i}$ possess. The winsorization/trimming used in the present article draws from ideas in Lugosi and Mendelson (2021) and Kock and Preinerstorfer (2025), is quantile-based, and “adapts” to $m$. Figure 1 illustrates, in particular, that the winsorized mean we study obeys a tight Gaussian approximation even for rather large $d$.
Furthermore, our results only require $\log(d)=o\mathinner{\bigl(n^{\frac{m-2}{5m-2}}\bigr)}$, which is exponential for all $m>2$, albeit with a small exponent for $m$ close to two. As $m\to\inftydd$ is allowed to grow almost as fast as $\exp(n^{1/5})$, improving on the rate in Resende (2024), and “recovering” the best known rate for Gaussian approximations based on the sample mean of $X_{i}$ with sub-exponential entries (cf. Remark 2 in Chernozhukov et al. (2023a)). This rate remains valid for the bootstrap procedures that we consider and all results are robust to some adversarial contamination.
Our bootstrap approximations are based on a novel “winsorized” covariance matrix estimator, for which we establish performance guarantees in the presence of adversarial contamination and only requiring $m>2$ moments. Because the winsorization mechanism used in the construction of that covariance estimator is quantile-based, the estimator does not require knowledge of any unknown population quantities. In particular, its performance guarantees adapt to the unknown $m$, which may be of independent interest.
Gaussian approximations for winsorized means are obtained in Section 2, whereas bootstrap approximations are studied in 3. Some results on trimmed means are established in Section 5. Numerical results in addition to the ones already discussed in Figure 1 above are presented in Section 6.
@@ 2 Gaussian approximations for winsorized means
Recall that $X_{1},\ldots,X_{n}$ is a sample of i.i.d. random vectors in $\mathbb{R}^{d}$, and where $X_{i}=(X_{i,1},\ldots,X_{i,d})^{\prime}$ for $i=1,\ldots,n$. Let $\mu=(\mu_{1},\ldots,\mu_{d})^{\prime}=\mathbb{E}X_{1}$, $\Sigma$ be the covariance matrix of $X_{1}$, and for $m\in[2,\infty)$ let $\sigma_{m,j}^{m}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}|X_{1,j}-\mu_{j}|^{m}$, all of which are well-defined under Assumption 2.1 below. We suppress the dependence of $d=d(n)$ on $n$ in our notation.
@@
where $\overline{\eta}_{n}\in(0,1/2)$ is a non-random and known upper bound on the fraction of contaminated observations. Which of $\tilde{X}_{i}$ differ from $X_{i}$ as well as their values can depend on the uncontaminated sample $X_{1},\ldots,X_{n}$. Adversarial contamination has become a popular criterion to evaluate robustness of a statistic against, as it allows for many forms of data manipulation, cf. Lai et al. (2016), Cheng et al. (2019), Diakonikolas et al. (2019), Hopkins et al. (2020), Lugosi and Mendelson (2021), Minsker and Ndaoud (2021), Bhatt et al. (2022), Depersin and Lecué (2022), Dalalyan and Minasyan (2022), Minasyan and Zhivotovskiy (2023), Minsker (2023), Oliveira et al. (2025). The recent book by Diakonikolas and Kane (2023) provides further references and discussion of various contamination settings. Since the sample mean has a breakdown point of $1/n$, Gaussian approximations based on $S_{n}$ are not robust to adversarial contamination (or large outliers).
@@
The $X_{1},\ldots,X_{n}$ are i.i.d. random vectors in $\mathbb{R}^{d}$ with $\mathbb{E}|X_{1,j}|^{m}<\infty$ for some $m\in(2,\infty)$ and all $j=1,\ldots,d$. There exist strictly positive constants $b_{1}$ and $b_{2}$ such that $\min_{j=1,\ldots,d}\sigma_{2,j}\geq b_{1}$ and $\sigma_{m}\mathrel{\mathop{\ordinarycolon}}=\max_{j=1,\ldots,d}\sigma_{m,j}\leq b_{2}$. The actually observed adversarially contaminated random vectors (in $\mathbb{R}^{d}$) are denoted $\tilde{X}_{1},\ldots,\tilde{X}_{n}$ and satisfy (2).
@@ 2.1 The winsorized means
with $\hat{\alpha}_{j}=\tilde{X}_{\lceil\varepsilon_{n}n\rceil,j}^{*}$ and $\hat{\beta}_{j}=\tilde{X}_{\lceil(1-\varepsilon_{n})n\rceil,j}^{*}$ for a carefully chosen $\varepsilon_{n}\in(0,1/2)$. Thus, for each coordinate $j$, the winsorization points $\hat{\alpha}_{j}$ and $\hat{\beta}_{j}$ are order statistics of the contaminated data $\tilde{X}_{1,j},\ldots,\tilde{X}_{n,j}$. In this sense, the winsorization mechanism employed is quantile-based. [^2] For the purpose of constructing estimators of $\mu\in\mathbb{R}$ with finite-sample sub-Gaussian concentration properties, related winsorized mean estimators were recently studied in Lugosi and Mendelson (2021) and Kock and Preinerstorfer (2025). The winsorized estimators studied in the latter article remove some practical limitations of the ones studied in the former. We therefore mainly build on the results in Kock and Preinerstorfer (2025). Under adversarial contamination it is clear that even $S_{n,W}$ can perform arbitrarily badly unless at least the smallest and largest $\overline{\eta}_{n}n$ observations are winsorized. Thus, one must choose $\varepsilon_{n}\geq\overline{\eta}_{n}$. Our Gaussian approximations below apply for any $\varepsilon_{n}\in(0,0.5)$ of the form
$\varepsilon_{n}=\lambda_{1}\cdot\overline{\eta}_{n}+\lambda_{2}\cdot\frac{\log(dn)}{n},\qquad\text{with }\lambda_{1}\in(1,\infty)\text{ and }\lambda_{2}\in(0,\infty).$ (4)
We impose $\lambda_{1}>1$ since otherwise even $S_{n,W}$ can be arbitrarily manipulated by the adversary. To make efficient use of the data, we recommend $\lambda_{1}$ close to but greater than one. In the leading special case of $\overline{\eta}_{n}=0$, which is the setting in which high-dimensional Gaussian approximations for the sample mean based $S_{n}$ have been studied (cf. the literature summarized in the introduction), one has that $\varepsilon_{n}=\lambda_{2}\cdot\log(dn)/n$. The simulations in Section 6 suggest that $\lambda_{2}\approx 0.1$ works well in practice.
Without further mentioning, $\lambda_{1}\in(1,\infty)$ and $\lambda_{2}\in(0,\infty)$ are taken as fixed throughout the remainder of the paper.